OpenStudy (anonymous):

Milk and cream are mixed together for a recipe. The total volume of the mixture is 1 cup. If the milk contains 2% fat, the cream contains 18% fat, and the mixture contains 6% fat, how much cream is in the mixture?

OpenStudy (danjs):

1 = M + C (2/100)* M + (18/100)*C = 6/100 2 equations, 2 unknowns

OpenStudy (danjs):

Is that what you are studying at the time? Systems of equations?

OpenStudy (anonymous):

yes

OpenStudy (danjs):

Ok, the hard part i am guessing, is translating the paragraph into the equations right?

OpenStudy (anonymous):

oh cool thx

OpenStudy (danjs):

Here..

OpenStudy (danjs):

Define the variables first... Let M = cups of milk C = cups of cream

OpenStudy (danjs):

The total volume of the mixture is 1 cup Total Mixture = Cups of Cream + Cups of Milk 1 Cup = C cups of cream + M cups of milk 1 = C + M

OpenStudy (danjs):

If the milk contains 2% fat, the cream contains 18% fat, and the mixture contains 6% fat (Fat From Milk)*(Cups of milk) + (fat From Cream)*(Cups of Cream) = Fat in mixture (0.02)*M + ( 0.18)*C = (0.06)

OpenStudy (danjs):

So the 2 equations to be solved are C + M = 1 (0.02)*M + ( 0.18)*C = (0.06) For C cups of cream, and M cups of milk..

OpenStudy (danjs):

you with me so far?

OpenStudy (danjs):

Solve first equation for either M or C, i will solve it for C. M + C = 1 C = 1 - M Substitute C = 1 - M into the second equation (0.02)M + 0.18*(1 - M) = 0.06

OpenStudy (danjs):

One variable in one equation, we can solve for the M cups of milk 0.02M + 0.18 - 0.18M = 0.06 -0.16M = -0.12 M = (-0.12)/(-0.16) M = 0.75 cups of milk

OpenStudy (danjs):

If M = 0.75 cups milk, then from the first equation, the total mixture is 1 cup M + C = 1 0.75 + C = 1 C = 1 - 0.75 C=0.25 cups Cream

OpenStudy (danjs):

The mixture contains 0.75 cups of Milk 0.25 cups of cream

OpenStudy (danjs):

@aeh101